If ABC
+ DEF
= GHI
and
ADG
+BEH
=CFI
What numbers do you plug into this equation to make it correct both ways. IT says there are two answers.

Answer by Edwin McCravy(20054)


There are 4 solutions if you DON'T allow using zeros.
Actually there really are only two pairs because one
is just the other one switched:
 
ABC + DEF = GHI     ADG + BEH = CFI
146 + 583 = 729     157 + 482 = 639

ABC + DEF = GHI     ADG + BEH = CFI
157 + 482 = 639     146 + 583 = 729

-----------------------------------

ABC + DEF = GHI     ADG + BEH = CFI
718 + 236 = 954     729 + 135 = 864

ABC + DEF = GHI     ADG + BEH = CFI
729 + 135 = 864     718 + 236 = 954

-----------------------------------

However, if you allow 0's for the second or third
digits, there are two more pairs of solutions:

ABC + DEF = GHI     ADG + BEH = CFI
326 + 584 = 910     359 + 281 = 640

ABC + DEF = GHI     ADG + BEH = CFI
359 + 281 = 640     326 + 584 = 910

-----------------------------------

ABC + DEF = GHI     ADG + BEH = CFI
348 + 562 = 910     359 + 461 = 820

ABC + DEF = GHI     ADG + BEH = CFI
359 + 461 = 820     348 + 562 = 910

-----------------------------------
Edwin